A family has two cars. During one particular week, the first car consumed
40
gallons of gas and the second consumed
30
gallons of gas. The two cars drove a combined total of
1900
miles, and the sum of their fuel efficiencies was
55
miles per gallon. What were the fuel efficiencies of each of the cars that week?

Answer by Boreal(15235)

40 gallons* x mpg=40x miles, so x mpg is the fuel efficiency of the first car
30 gallons*(55-x)mpg=1650-30x miles
That sum is 1900 miles
so 10x+1650=1900
10x=250
x=25 mpg for car 1 and went 1000 miles
55-x=30 mpg for car 2 and went 900 miles